OpenStudy (anonymous):

a function f is an even function if f(-x)=f(x) for all x and is an odd function if f(-x)=-f(x) for all x. Prove that the derivative of an even function is odd and the derivative of an odd function is even

OpenStudy (hunus):

even \[f(-x)=f(x)\] odd \[f(-x)=-f(x)\] \[f(x)=a_{n}x^n+a_{n-1}x^{n-1}+a_{n-2}x^{n-2}+...+a_{1}x+a_{0}\] for even polynomial functions n must be even for odd polynomial functions n must be odd Given an even function of the form \[f(x)=a_{n}x^n+a_{n-1}x^{n-1}+a_{n-2}x^{n-2}+...+a_{1}x+a_{0}\] \[f'(x)=na_{n}x^{n-1}+(n-1)a_{n-1}x^{n-2}+(n-2)a_{n-2}x^{n-3}+...+a_{1}\] Lowering the power of the even function \[n \rightarrow (n-1)\] by one results in an odd function etc etc